### 17 Dec 2023 v1

# AI Gender Bias, Disparities, and Fairness: Does Training Data Matter?

Abstract.
This study delves into the pervasive issue of gender issues in artificial intelligence (AI), specifically within automatic scoring systems for student-written responses. The primary objective is to investigate the presence of gender biases, disparities, and fairness in generally targeted training samples with mixed-gender datasets in AI scoring outcomes. Utilizing a fine-tuned version of BERT and GPT-3.5, this research analyzes more than 1000 human-graded student responses from male and female participants across six assessment items. The study employs three distinct techniques for bias analysis: Scoring accuracy difference to evaluate bias, mean score gaps by gender (MSG) to evaluate disparity, and Equalized Odds (EO) to evaluate fairness. The results indicate that scoring accuracy for mixed-trained models shows an insignificant difference from either male- or female-trained models, suggesting no significant scoring bias. Consistently with both BERT and GPT-3.5, we found that mixed-trained models generated fewer MSG and non-disparate predictions compared to humans. In contrast, compared to humans, gender-specifically trained models yielded larger MSG, indicating that unbalanced training data may create algorithmic models to enlarge gender disparities. The EO analysis suggests that mixed-trained models generated more fairness outcomes compared with gender-specifically trained models. Collectively, the findings suggest that gender-unbalanced data do not necessarily generate scoring bias but can enlarge gender disparities and reduce scoring fairness.
Artificial Intelligence (AI), Gender Bias, Machine Learning, Automatic Scoring, Education, Equalized Odds
Figure 1. Overview of AI gender bias analysis for automatic scoring.
1. Introduction
Artificial Intelligence (AI) has permeated various aspects of human life, revolutionizing industries and practices with its advanced computational capabilities. However, the potential of AI is often marred by inherent biases, especially gender biases, that can have profound implications on the fairness and effectiveness of AI applications. Gender bias in AI, particularly in the context of automatic scoring systems, poses a significant challenge, echoing societal stereotypes and disparities (Zhai and Nehm, 2023). Studies such as (Leavy, 2018; Nadeem et al., 2020; Cirillo et al., 2020) have highlighted the critical need for diversity and gender theory in machine learning to address these biases.
Pseudo-AI bias, a concept introduced by Zhai and Krajcik (Zhai and Krajcik, 2022), counters the mechanical bias inherent in AI systems and often draws unnecessary anxiety about AI. The mechanical bias, while not intentional, can lead to discriminatory outcomes, especially in gendered contexts. The work of Bolukbasi et al. (Bolukbasi et al., 2016) and Hall et al. (Hall and Ellis, 2023) further underscores the impact of such biases stemming from word embeddings and supervised machine learning models, respectively.
The significance of this study lies in its focus on providing statistical evidence regarding the gender-neutral nature of the mixed-gender trained model by analyzing more than 1000 human-graded student responses across different gender categories. We have trained models using male-only, female-only, and mixed written datasets and evaluated the performance of these models against males and females only and mixed testing data. This research employs standard techniques to measure gender bias using scoring accuracy difference (pair t-test (Mowery, 2011)), gender disparity using mean score gaps between predicted scores for different genders (Wang and Brown, 2007; Wilson et al., 2023), and fairness using Equalized Odds (Hardt et al., 2016), to evaluate gender bias, discrepancies, and fairness (See Fig 1 for a detailed overview of the approach). The findings reveal a promising trajectory toward analyzing gender issues, evidenced by the insignificant accuracy difference that means non-biased prediction, low MSGs mean non-disparate prediction, and minimal odds gradient gamma (¡0.01) mean fair prediction for mixed and separate models.
This study contributes to the ongoing discourse on gender bias in AI, aligning with recent research (Franzoni, 2023; Lima et al., 2023; Hall and Ellis, 2023) that emphasizes the need for gender-sensitive AI development. Furthermore, it resonates with the efforts in mitigating gender bias through comprehensive reviews and algorithmic adjustments (Sun et al., 2019; Li et al., 2022; Lu et al., 2020).
The motivation behind this study is rooted in the ethical imperative to ensure fairness and equity in AI systems, a sentiment echoed by Slimi (Slimi, 2023) and O’Connor & Liu (O’Connor and Liu, 2023). By demonstrating that targeted mix model training doesn’t have traces of gender bias, this research offers a blueprint for developing more equitable AI systems in education and beyond, contributing to the broader discourse on ethical AI (Nemani et al., 2023; Manresa-Yee and Ramis, 2021).
2. Background
2.1. AI in Education
Integrating AI in education marks a paradigm shift in teaching and learning. AI’s role in education encompasses a broad spectrum, from personalized learning to administrative efficiency and pedagogical innovations.
Roles and Impacts: AI’s implementation in educational contexts is multifaceted. Studies (Holmes et al., 2021; Berendt et al., 2020) highlight the ethical dimensions of AI in education, emphasizing the need for frameworks that balance technological advancement with moral and social values. Felix (Felix, 2020) and Qadir (Qadir, 2023) discuss the evolving role of AI, from augmenting teaching methodologies to supporting administrative tasks, offering insights into how AI can enhance the educational experience for both teachers and students.
Educational Analytics and Assessment: AI-driven analytics play a crucial role in assessing student performance and identifying learning gaps. By processing large volumes of data, AI systems can provide educators insights into student progress, enabling more informed decision-making and targeted instructional strategies.
Ethical and Social Considerations: The incorporation of AI in education also raises important ethical questions. Data privacy, algorithmic fairness, and the digital divide are central to the discourse on AI in education. A recent study (Holmes et al., 2021) underscores the importance of addressing these ethical concerns, advocating for a community-wide framework that ensures AI is used responsibly and equitably in educational settings.
AI’s role in education rapidly evolves, offering exciting possibilities for enhancing teaching and learning experiences. However, the successful integration of AI in education hinges on addressing the ethical, social, and practical challenges accompanying these technological advancements.
2.2. Automatic Scoring in Education
The advent of AI in education has brought significant advancements in automatic scoring systems, revolutionizing how student assessments are conducted and analyzed (Zhai et al., 2020b, a; Latif and Zhai, 2023a). This technology offers scalability and efficiency for instruction and introduces new paradigms in educational assessment practices.
Current State and Advancements: González-Calatayud et al. (González-Calatayud et al., 2021) provide a comprehensive systematic review of AI applications in student assessment, highlighting various methodologies and their effectiveness. Zhai et al. (Zhai et al., 2021a) reviewed AI-based assessments in science and discuss the state of the art and practice, focusing on the latest advancements and research findings. Their work illustrates how AI technologies, particularly machine learning and natural language processing, have effectively evaluated student responses in diverse subjects, though with varying degrees of scoring accuracies.
Challenges in Implementation: Despite these advancements, implementing automatic scoring is challenging. Studies (Zhai et al., 2022) delve into the application of machine learning in scientific assessments, shedding light on the complexities and nuances involved in accurately assessing student models. Zhai & Nehm (Zhai and Nehm, 2023) further explore these challenges, particularly in formative assessments, and discuss how AI can be optimized for more effective educational outcomes.
Machine-Human Score Agreements: A critical aspect of automatic scoring is the agreement between machine-generated scores and human evaluations. A meta-analysis of machine learning-based science assessments was conducted to understand the factors impacting machine-human score agreements (Zhai et al., 2021b). Their findings suggest that while AI can closely mimic human scoring in certain contexts, discrepancies still exist, underscoring the need for continuous refinement of AI algorithms in educational settings.
Potential and Future Directions: The potential of AI in automatic scoring extends beyond efficiency and scalability. It offers opportunities for personalized feedback, adaptive learning environments, and a deeper analysis of student understanding (Zhai, 2021; Lee et al., 2023b). As suggested by researchers, the future of AI in education lies in generative AI’s ability to complement and enhance traditional teaching methods, providing valuable insights into student learning processes and outcomes. LLMs represent a transformative tool with immense potential. Researchers have leveraged these models in automatic scoring of science writing (Latif and Zhai, 2023b; Lee et al., 2023a).
2.3. AI Gender Bias, Disparities, and Fairness
Gender bias, disparities, and fairness in AI, particularly in education, is a topic shrouded in both myths and harsh realities. We consider the systematic definition of bias for AI (similar to (Larrazabal et al., 2020)), which refers to systematic rather than random error in which we analyze the calculated error in scoring accuracy of trained AI models. Gaps in automatically predicted machine scores for responses written by different genders of students are considered gender disparities in AI (concept defined by OECD (for Economic Co-operation and, OECD)). Further, AI fairness as equity (Holstein and Doroudi, 2019) is considered where the AI model assumes to be fair if it demonstrates an equal proportion of false positives and true positives for automatically predicted scores of different gender written responses. This section aims to dissect these aspects, providing a clearer understanding of the extent and implications of gender issues in AI systems.
Unraveling Myths: Common myths around AI gender bias often stem from misunderstandings of AI functionalities and oversimplified interpretations of gender disparities. One such myth is the belief that AI systems are inherently unbiased as they are based on algorithms and data. However, as Bolukbasi et al. (Bolukbasi et al., 2016) and Hardt et al. (Hardt et al., 2016) illustrate, biases in training data and algorithmic design can lead to gender-biased outcomes. Another myth is that gender bias in AI is solely a technical issue that can be resolved with more data or better algorithms. While these are crucial factors, as pointed out by Lu et al. (Lu et al., 2020) and Manresa-Yee & Ramis (Manresa-Yee and Ramis, 2021), addressing gender bias in AI also requires a broader understanding of societal and cultural contexts.
Confronting Realities: The reality of gender bias in AI is multifaceted and deeply rooted in both the technical and social domains. The concept of pseudo-AI bias (Zhai and Krajcik, 2022) highlights how biases can be subtly embedded in AI systems, often reflecting societal stereotypes. These biases can manifest in various forms, such as skewed representations in datasets or biased interpretations of AI outcomes.
Impact on Education: In educational settings, AI gender bias can have significant consequences. It can lead to unfair assessment practices, reinforce gender stereotypes, and create barriers to equal educational opportunities. Madaio et al. (Madaio et al., 2022) delve into the structural injustices that AI can perpetuate in educational contexts, advocating for a more holistic approach to understanding and addressing these biases.
Debunking Myths and Embracing Realities: Ultimately, debunking myths and embracing the realities of AI gender bias in education is crucial for developing AI systems that are fair, equitable, and beneficial for all learners. This calls for a multi-disciplinary approach incorporating insights from gender studies, ethical AI practices, and educational theory to ensure that AI is a tool for empowerment rather than perpetuating existing disparities.
Understanding and addressing AI gender bias in education requires moving beyond myths and confronting the complex realities of how gender biases are embedded and perpetuated in AI systems. This understanding is crucial for developing AI applications that are equitable and effective in diverse educational contexts.
Study’s Approach to Bias, Disparities, and Fairness Analysis: The study employs a rigorous methodology to analyze gender issues in automatic scoring. Utilizing a fine-tuned version of BERT (similar to (Liu et al., 2023)) and GPT-3.5 (similar to (Latif and Zhai, 2023b)), the research analyzes around 1000 human-graded student responses, proportionally distributed among male, female, and mixed datasets across six assessment items. The study measured bias by pair t-test of scoring accuracy between different gender-specifically trained models (Mowery, 2011); the gender disparities by mean score gaps by gender(Wang and Brown, 2007; Wilson et al., 2023); and fairness by EO (Hardt et al., 2016). These methods comprehensively show how gender issues manifest in AI scoring systems for mixed and separately trained models.
Gender-Neutral Nature of Mixed Trained Models: The core hypothesis of the study is that training AI models with mixed-gender datasets doesn’t show gender preferences for automatic scoring upon testing male and female responses. This hypothesis is grounded in the understanding that diverse training data can help AI algorithms learn more nuanced, gender-neutral patterns, thus mitigating biases that arise from unbalanced or skewed datasets.
Setting Ground to Avoid AI gender issues: A critical aspect of this study is its commitment to dispelling myths surrounding AI and gender issues. By employing a data-driven, empirical approach, the study demonstrates that gender issues in AI are not immutable and can be addressed through thoughtful and deliberate training methods. This counters the common misconception that AI systems are inherently biased or that biases in AI are too complex to be resolved.
Implications for AI Development: The findings of this study have significant implications for AI development, particularly in educational contexts. They suggest carefully considering training data and applying targeted analysis techniques can lead to more equitable and fair AI systems. This approach aligns with the broader discourse on ethical AI development, emphasizing the need to incorporate diversity and gender theory in AI training practices.
To sum up, this study contributes to the ongoing efforts to analyze many gender issues in AI through targeted training and rigorous analysis. By providing empirical evidence and a methodological framework, it sets the ground for developing AI systems that are both technically sound and ethically responsible, thereby avoiding common myths.
3. Methodology
This study aims to investigate gender issues in AI-based scoring systems based on State-Of-The-Art Language Models BERT and GPT-3.5 models. Large Language models tend to revolutionize education technology (Latif et al., 2023) and raise potential challenges such as their bias, disparity, and fairness, which need to be addressed. A comprehensive methodology was adopted, comprising different training and testing phases for models and employing various statistical techniques to evaluate gender bias, disparities, and fairness.
3.1. Dataset
The study utilized a meticulously categorized dataset of student-written responses to six assessment items, each classified under the multi-class category. The datasets were categorized based on reported gender (i.e., male and female; we eliminated the data that did not report gender information), and a mixed dataset incorporated responses from both genders. Based on the availability of data, we made use of the maximum of available data while maintaining the ratio between training and testing data for each item by gender. Consequently, we randomly identified three training datasets for each item: mixed, male, and female training datasets; and three testing datasets for each item. The detailed composition of each assessment item’s dataset is presented in Table 1.
Dataset Overview:
  • Falling Weights: Includes 1148 training and 230 testing samples in the mixed dataset, with 342 training and 87 testing samples for Males and 478 training and 120 testing samples for females.
  • Gelatin: Comprises 918 training and 230 testing samples in the mixed dataset, with 344 training and 86 testing samples for Males and 477 training and 120 testing samples for females.
  • Bathtub: Contains 916 training and 230 testing samples in the mixed dataset, with 364 training and 92 testing samples for Males and 450 training and 113 testing samples for females.
  • Sandwater1 and Sandwater2: Both have similar structures, with Sandwater1 having 915 training and 229 testing samples and Sandwater2 having 913 training and 229 testing samples in the mixed dataset. For both, Males contributed 364 training and 92 testing samples, while Females contributed around 450 training and 113 testing samples.
  • Two Boiling Situations: Consists of 909 training and 228 testing samples in the mixed dataset, with 363 training and 91 testing samples for Males and 448 training and 113 testing samples for females.
This comprehensive dataset facilitated a nuanced analysis of gender bias in AI-based scoring systems, ensuring robust and broadly applicable study findings.
Table 1. Summary of Datasets and Their Samples
Datasets Category mixed Training samples mixed Testing samples Male Training Male Testing Female Training Female Testing
falling weights
Multi-class
1148
230
342
87
478
120
gelatin
Multi-class
918
230
344
86
477
120
bathtub
Multi-class
916
230
364
92
450
113
sandwater1
Multi-class
915
229
364
92
450
113
sandwater2
Multi-class
913
229
364
92
449
113
two boiling situations
Multi-class
909
228
363
91
448
113
3.2. Experimental Setup
The experimental setup involved rigorous training and testing using multiple training and testing datasets.
Training Phase: Each dataset undergoes a unique training cycle:
  • A mixed-trained model trained with the mixed training dataset including both genders.
  • A male-trained model trained exclusively with data from males.
  • A female-trained model trained solely with data from females.
Using this approach, we developed three distinct language models for each assessment item, fine-tuned using both BERT and GPT-3.5, respectively.
Testing Phase: Each trained model undergoes multiple testing phases to detect potential biases. Specifically,
  • Experiment 1. The model trained on the mixed dataset was tested against mixed testing data and separately against testing data from males and females. We used both BERT and GPT-3.5.
  • Experiment 2. Models trained on a single gender’s data were tested on their respective gender’s test data and against the test data from the opposite gender. Similarly, we used both BERT and GPT-3.5.
This experimental design aimed to rigorously assess any inherent biases in the models’ predictions across different datasets and gender divisions.
3.3. Accuracy Difference (Pair t-Test)
The accuracy difference between models was assessed using a paired t-test, a statistical method ideal for comparing the means of two related groups. Mowery (Mowery, 2011) describes the paired t-test as a powerful tool for detecting differences in paired observations. Our study involved comparing the accuracies of models trained on different datasets.
Mathematically, the paired t-test is expressed as:
(1) $t=\frac{\overline{D}}{s_{D}/\sqrt{n}}$
where $\overline{D}$ is the mean of the differences between paired observations, $s_{D}$ is the standard deviation of the differences, and $n$ is the number of pairs. A p-value greater than 0.05 indicated no significant difference in accuracies between the models, suggesting minimal gender bias.
3.4. Mean Score Gap
Mean Score Gap (MSG) was utilized to evaluate the disparity in scores between different genders. MSG can be generated either by human experts or machine algorithmic models. Given that human-assigned scores have undergone rigorous scrutiny, we used MSG generated by human-assigned scores as the standard to evaluate machine-generated MSG. It can be seen that if machine-generated MSG is larger than human-generated MSG for the same testing data, the machine exacerbates the gender difference and may potentially create inequity. Wang and Brown (Wang and Brown, 2007) and Wilson et al. (Wilson et al., 2023)highlighted the importance of such comparisons in their study on automated essay scoring and automatic scoring of written scientific arguments. MSG is calculated as:
(2) $MSG=\frac{1}{n}\sum_{i=1}^{n}|score_{male,i}-score_{female,i}|$
where $score_{gender1,i}$ and $score_{gender2,i}$ are the scores assigned to the $i^{th}$ response by models trained on Male and Female datasets, respectively, and $n$ is the total number of responses. A threshold of MSG ¡ 0.2 was used to indicate acceptable score disparity.
3.5. Equalized Odds
Equalized Odds, as defined by Hardt et al. (Hardt et al., 2016), involves assessing the equality of true and false positive rates across groups. It is a crucial measure of fairness in predictive models. The Equalized Odds (EO) criterion is given by:
(3) $EO=max(|TPR_{male}-TPR_{female}|,|FPR_{male}-FPR_{female}|)$
where $TPR_{male}$ and $TPR_{female}$ are the true positive rates for Male and Female, respectively, and $FPR_{male}$ and $FPR_{female}$ are the false positive rates for the two genders. An EO value less than 0.01 indicates a fair model with minimal gender bias.
These three methods - Paired t-test, MSG, and Equalized Odds - provide a comprehensive approach to evaluating gender bias, disparities, and fairness in AI-based scoring systems, ensuring the ethicaal uses of AI models. The methodology encompassed a detailed experimental setup with rigorous training and testing phases, coupled with statistical analyses to assess gender issues in AI-based scoring systems comprehensively.
4. Results
4.1. Scoring Accuracy Difference Evaluation
A paired samples t-test was performed to evaluate whether there was a statistical difference between the scoring accuracy for male and female predicted scores before and after randomly mixing both genders’ responses.
4.1.1. BERT Model Results
[…]
4.1.2. GPT-3.5 Turbo Model Results
[…]
Interpretation: Both the BERT and GPT-3.5 models demonstrated consistent performance across mixed and gender-specific datasets, suggesting minimal gender biases.
Figure 2. Mean score gaps between male and female testing data: Comparing human-graded scores and fine-tuned LLMs’ scores: Mixed training Model (Left), Male trained Model (Center), and Female trained Model (Right) comparison line plots of BERT (Top) and GPT3.5 (Bottom) fine-tuned models’s scores.
4.2. Mean Score Gap
To verify our assumption, we designed two experiments to compare the difference of MSGs between human- and machine-graded scores. We tested the three models: Mixed-, male-, and female-trained models, and compared each of them with human-generated MSG.
4.2.1. Experiment 1: Mean Score Gap for the Mixed-Training Models
The first experiment analyzed the MSGs between human and mix-trained machine-graded scores for BERT and GPT-3.5, respectively. We tested the model performance with mixed, male, and female testing datasets, respectively.
For the BERT model, the MSG of the six tasks yielded from human-assigned scores on the mixed testing dataset was $M_{mixed}=0.032$, the male testing data was $M_{male}=0.056$, and the female testing data was $M_{female}=0.042$. Correspondingly, MSG generated by the mixed-trained BERT model on the three testing datasets were $M_{mixed}=0.022$, $M_{male}=0.048$, and $M_{female}=0.033$. The results indicate that the MSG between mixed test datasets against humans ($MD_{mixed}=0.011$) is comparatively lower than that of males ($MD_{male}=0.08$) and females ($MD_{female}=0.09$).
In contrast, for the GPT-3.5 model, the MSG of the six tasks yielded from human-assigned scores on the mixed testing dataset was $M_{mixed}=0.074$, the male testing data was $M_{male}=0.082$, and the female testing data was $M_{female}=0.112$. Correspondingly, MSG generated by the mixed-trained GPT-3.5 model on the three testing datasets were $M_{mixed}=0.072$, $M_{m}ale=0.072$, and $M_{female}=0.105$. The results indicate that the MSG between mixed test datasets against humans ($MD_{mixed}=0.002$) is comparatively lower than that of males ($MD_{male}=0.010$) and females ($MD_{female}=0.07$).
4.2.2. Experiment 2: Mean Score Gap for Gender Specifically Trained Models
To understand the MSGs generated by the gender-specifically-trained models, the second experiment compared the MSGs yielded by machine-graded scores and human-graded scores, respectively. For each assessment task, we calculated MSG on the male and female testing datasets, respectively. We first reported the BERT results, followed by the GPT-3.5 results.
For the BERT model, overall the two gender-specifically trained models demonstrated a larger average MSG for the six tasks, M = 0.107, compared to that generated by the human grader, M = 0.98. For example, ’sandwater2’, the MSGs were close for both testing datasets (Males: M=2.38, SD=0.67; Females: M=2.20, SD=0.73), akin to human MSGs (Males: M=2.33, SD=0.71; Females: M=2.20, SD=0.74), respectively. When a gender-specifically trained male model was tested on gender-counterpart testing data, we found that both Male and Female-trained Models showed increased MSGs on both testing datasets compared to human-generated MSGs. For example, the Male model in ’twoboilingsituation’ scored Female testing data higher (M=2.42, SD=0.59) than Female testing data, indicating an increased gender disparities. Conversely, when gender-specific models were tested on the same gender testing datasets, they exhibited decreased gender disparities (Detailed results can be seen in Fig. 2). An anomaly noted was the similarity in standard deviations between the Mixed Model and human scores, indicating consistent scoring variance.
Similarly, the GPT-3.5, overall, the fine-tuned male- and female-trained models exhibited a larger average MSG for six tasks, M = 0.115, compared to that graded by the human expert, M = 0.98. For instance, for the item ’twoboilingsituation’, the MSGs were close for both testing datasets (Males: M=2.40, SD=0.88; Females: M=2.15, SD=0.83), mirroring human scores (Males: M=2.37, SD=0.96; Females: M=2.19, SD=0.87), respectively. When a gender-specifically trained model was tested on gender-counter part testing data, we found that both Male and female-trained models showed increased MSGs on both testing datasets compared to human-generated MSGs. For instance, the male Model in ’sandwater1’ had a lower score for females (M=2.12, SD=0.76) than males. Across all items, the Mixed testing’s scores closely resembled human graders, suggesting reduced bias (Detailed results can be seen in Fig. 2).
Interpretation: These results indicate that training with a mixed dataset in both BERT and GPT-3.5 models has shown a minimal MSG compared to gender-specific training. This outcome validates our hypothesis that a balanced mixed training dataset have significant traces toward gender disparitiess, as reflected in the reduced MSGs in the mixed models. The findings underscore the importance of inclusive training datasets in developing AI models that yield more equitable scoring outcomes.
4.3. Equalized Odds Evaluation
The study evaluated the Equalized Odds (EO) (Hardt et al., 2016) for BERT and GPT-3.5 models to assess fairness in gender predictions.
4.3.1. BERT Model Equalized Odds
The BERT model demonstrated an EO of 0.042 on the mixed testing dataset for the mixed-trained model, indicating a lower disparity in prediction fairness. In contrast, the gender-specifically-trained models exhibited higher EO values on the mixed testing dataset, with the Male model showing an EO of 0.107 and the Female model having an EO of 0.074. These results suggest that the mixed-trained model is more equitable in its predictions across genders than gender-specific models.
4.3.2. GPT-3.5 Model Equalized Odds
For the GPT-3.5 model, the mixed-trained model’s EO was 0.061 on the mixed testing dataset, again reflecting a reduced gradient and, thus, greater fairness in predictions. The Male and Female trained models showed EOs on the same testing dataset of 0.076 and 0.074, respectively. While these values are closer to that of the mixed model than the BERT model, they indicate a higher disparity in for the gender-specific models.
Interpretation: The mixed trained models for both BERT and GPT-3.5 have lower EO values, suggesting more equitable predictions and higher fairness compared to gender-specific models. This outcome supports the hypothesis that training with balanced dataset can reduce the EO gradient, contributing to the development of fair AI models.
5. Discussion
This study demonstrates that gender unbalanced training datasets doesn’t have traces of gender bias for AI-based scoring systems. The prevalent notion that AI inherently perpetuates gender biases (Bolukbasi et al. (Bolukbasi et al., 2016); Lu et al. (Lu et al., 2020)) is challenged by the outcomes of this study. However, our research shows that models trained on data lacking gender balance might lead to algorithms exacerbating gender disparities. The reduced MSG and lower EO gradients in models trained on mixed datasets highlight the potential of balanced training in addressing gender fairness. This finding is crucial in the context of Hardt et al. (Hardt et al., 2016), who emphasize the importance of equality of opportunity in supervised learning.
In education, where AI’s fairness is paramount (Li et al. (Li et al., 2022)), our study suggests that balanced mixed training datasets have the potential to enhance the equity of AI-based assessments. This approach aligns with the ethical frameworks proposed by Slimi (Slimi, 2023) for AI in higher education, emphasizing the need for inclusive and equitable AI practices.
The concept of pseudo AI bias (Zhai & Krajcik, 2022(Zhai and Krajcik, 2022)) implies that biases in AI are not always overt but can be subtle and deeply ingrained. The evidence of gender neutral behavior of the balanced mixed training model in this study provides a roadmap for addressing such false claims regarding biases in AI. By ensuring diverse representation in training data, AI models can be steered away from inadvertently reinforcing gender stereotypes.
Our findings resonate with the literature advocating for debiasing AI systems (Nadeem et al., 2022(Nadeem et al., 2022); Manresa-Yee & Ramis, 2021(Manresa-Yee and Ramis, 2021)). This study contributes to the broader efforts in creating fair and unbiased AI models by demonstrating that balanced training has more equitable AI outcomes. The approach counters the notion that AI systems are destined to replicate societal biases, offering a practical solution to this pervasive issue.
The results of this study present a compelling argument for the gender issue of balanced mixed training in AI for automatic scoring. This approach addresses the immediate concerns of fairness and equity in AI-based scoring systems and provides evidence against mechanical pseudo-bias in broader AI applications.
6. Conclusion
The analysis using BERT and GPT-3.5 on student written responses has significantly advanced our understanding of gender bias, disparities, and fairness in AI-based scoring systems. This study demonstrates that gender-unbalanced training has insignificant traces of gender bias. The findings challenge prevailing narratives about the inevitability of AI bias. We found that models trained on mixed-gender data, as opposed to gender-specific data, produced lower MSG when compared with human-generated results, implying that training data lacking gender balance might lead to algorithms exacerbating gender disparities. The EO analysis revealed that models trained on mixed-gender data achieved fairer outcomes than gender-specific ones. The results indicate that while gender-imbalanced data may not inherently lead to scoring bias, they can amplify gender disparities and diminish scoring fairness.
Acknowledgment
This study secondary analyzed data from project supported by the Institute of Education Sciences (grant number R305A160219). The authors acknowledge the funding agencies and the project teams for making the data available for analysis. Project is partially supported by NSF (grants numbers 2101104). The findings, conclusions, or opinions herein represent the views of the authors and do not necessarily represent the views of personnel affiliated with the funding agencies.