### 26 Dec 2023 v2

# AI Gender Bias, Disparities, and Fairness: Does Training Data Matter?

Abstract.
This study delves into the pervasive issue of gender issues in artificial intelligence (AI), specifically within automatic scoring systems for student-written responses. The primary objective is to investigate the presence of gender biases, disparities, and fairness in generally targeted training samples with mixed-gender datasets in AI scoring outcomes. Utilizing a fine-tuned version of BERT and GPT-3.5, this research analyzes more than 1000 human-graded student responses from male and female participants across six assessment items. The study employs three distinct techniques for bias analysis: Scoring accuracy difference to evaluate bias, mean score gaps by gender (MSG) to evaluate disparity, and Equalized Odds (EO) to evaluate fairness. The results indicate that scoring accuracy for mixed-trained models shows an insignificant difference from either male- or female-trained models, suggesting no significant scoring bias. Consistently with both BERT and GPT-3.5, we found that mixed-trained models generated fewer MSG and non-disparate predictions compared to humans. In contrast, compared to humans, gender-specifically trained models yielded larger MSG, indicating that unbalanced training data may create algorithmic models to enlarge gender disparities. The EO analysis suggests that mixed-trained models generated more fairness outcomes compared with gender-specifically trained models. Collectively, the findings suggest that gender-unbalanced data do not necessarily generate scoring bias but can enlarge gender disparities and reduce scoring fairness.
Artificial Intelligence (AI), Gender Bias, Machine Learning, Automatic Scoring, Education, Equalized Odds
Figure 1. Overview of AI gender bias analysis for automatic scoring.
1. Introduction
Today’s AI is so advanced in computation that it has penetrated deep into the well of human life, shaking up industries and practices. But the possibility of artificial intelligence is always spoiled by built-in biases, especially gender biases, which often cause serious problems. In particular, discrimination against women in AI takes the form of influential automatic scoring systems that reproduce societal stereotypes and gaps (Zhai and Nehm, 2023). Among others, see (Cirillo et al., 2020; Leavy, 2018; Nadeem et al., 2020) have stressed the important role that diversity is certain to play in machine learning over these questions of bias.
Zhai and Krajcik (2022) introduce the concept of pseudo-AI bias to combat this mechanical AI bias, often giving rise to unnecessary fear about artificial intelligence. Although unintentional, the mechanical bias leads to discriminatory outcomes. These become especially glaring where gender is involved. Such biases in exceedingly widely used word embeddings and supervised machine learning models are further underscored by the work of Bolukbasi et al. (2016) and Hall and Ellis (2023).
This is the significance of this study. By analyzing over 100 human-graded questions across various gender categories, it uses statistical evidence to show that both mixed-trained models are gender-neutral. Based on our male-only, female-only, and mixed written datasets (training), we have tested the performance of these models, creating both males and females or combined testing data. We use standard methods such as scoring accuracy difference (pair t-test (Mowery, 2011)) to measure potential gender bias, mean score gaps (Wang and Brown, 2007; Wilson et al., 2023)between machine and human scores for different genders to measure gender disparity, and Equalized Odds (Hardt et al., 2016) to measure fairness. (See Figure 1 for a detailed overview of the approach). The findings reveal a promising trajectory toward analyzing gender issues, evidenced by the insignificant accuracy difference that means non-biased prediction, low MSGs mean non-disparate prediction, and minimal odds gradient gamma (<.$<.01$) mean fair prediction for mixed and separate models.
Thus, this study adds to the current debate over gender bias in AI. It chimes with recent studies (Franzoni, 2023; Hall and Ellis, 2023; Lima et al., 2023) insisting on developing gender-aware artificial intelligence. Moreover, it is in tune with attempts to remove gender bias through all-round reviews and refinements of algorithms (Li et al., 2022; Lu et al., 2020; Sun et al., 2019).
According to Slimi (2023) and O’Connor and Liu (2023), this study is due to an ethical mandate of fairness or even equity in AI systems. This research shows that targeted mix model training lacks traces of gender bias, providing a pathway for creating more equitable AI in education and other contexts. It contributes to the growing debate about ethical or socially beneficial artificial intelligence AI (Manresa-Yee and Ramis, 2021; Nemani et al., 2023).
2. Background
2.1. AI in Education
AI, as an integral part of education, represents a fundamental change in the teaching and learning process. From personalized learning to administrative efficiency and pedagogical innovations, the fields in which AI can play a role are very diverse.
Roles and Impacts: Multifaceted implementation of AI in educational contexts. There are also studies (Holmes et al., 2021; Berendt et al., 2020) that discuss the ethical risks of AI in education and insist on keeping pace with technological progress while not violating moral or social values. How does AI fit into the educational process? Felix (2020) and Qadir Qadir (2023) share their perspectives on this, with changing applications of artificial intelligence from fortifying teaching methodologies to assisting administrative work.
Educational Analytics and Assessment: AI-based analysis is vital for evaluating student performance and discovering learning deficiencies. AI systems capable of handling large amounts of data can help educators interpret how students are doing so that they make more informed decisions and plan better-targeted approaches.
Ethical and Social Considerations: AI in education also raises important ethical questions. Algorithmic fairness, data privacy, and the digital divide are key to discussing AI in education. One recent study (Holmes et al., 2021) shares these ethical concerns, urging a community-wide framework to ensure that AI is used responsibly and equitably in educational settings.
The educational use of AI is developing rapidly, bringing new promise for enriching pedagogy and learning. Yet achieving the integration of AI in education depends on meeting the ethical, social and practical challenges posed by these technological developments.
2.2. Automatic Scoring in Education
Now that the era of AI in education is arriving, automatic scoring systems have evolved by leaps and bounds (Zhai et al., 2020b, a; Latif and Zhai, 2023a). Today’s student assessments differ greatly from those of yesteryear. It provides scalability and efficiency for training while promoting new paradigms in educational tests.
Current State and Advancements: In a recent and detailed systematic review of studies on AI applications in creating student assessments, González-Calatayud et al. (2021) cover methodologies that have been used with information about their effectiveness from different countries around the world in October 2015, Zhai et al. (2021a) had published a review of AI-based assessments on science that covered the state-of-the-art and practice concerning their latest developments as well as research outcomes. Their work shows that various AI technologies, especially machine learning and natural language processing, have already been proven to evaluate student answers in different subjects accurately.
Challenges in Implementation: These developments notwithstanding, making automatic scoring a reality is difficult. Zhai et al. (2022) explore the use of machine learning in scientific assessments, providing insight into how difficult it is to evaluate student models accurately. Zhai and Nehm (2023) look further into these problems, especially as they pertain to classroom assessments of student performance. They also consider how AI can be designed better for improved educational outcomes.
Machine-Human Score Agreements: Agreeing scores between machines and men is an important factor in automatic scoring. Zhai et al. (2021b) conducted a meta-analysis of machine learning-based science assessments to learn the reasons for low agreements between machine and human scores. Their conclusions indicate that although, in some aspects, AI can come extremely close to the humanly scored version, differences nonetheless remain. There is still a need to review and improve individual algorithm parameters within educational environments continually.
Potential and Future Directions: Improving efficiency and scalability is not all AI can do for automatic scoring. It provides room for personal feedback, an adapting learning environment, and an in-depth assessment of student comprehension (Lee et al., 2023b; Zhai, 2021). According to researchers, the future of AI in education is for generative AI technology to complement and augment verbal teaching methods. It could go further than that by even allowing us glimpses into student learning processes and outcomes like never before possible. With great potential, LLMs are a revolutionary tool. These models have been used by researchers in grading science writing automatically (Latif and Zhai, 2023b; Lee et al., 2023a).
2.3. AI Gender Bias, Disparities, and Fairness
Gender bias, disparities, and fairness in AI–especially educationally speaking, is a subject clouded by half-truths and hard facts. We consider the systematic definition of bias for AI (similar to (Larrazabal et al., 2020)), which refers to systematic rather than random error in which we analyze the calculated error in scoring accuracy of trained AI models. Gaps in automatically predicted machine scores for responses written by different genders of students are considered gender disparities in AI (concept defined by OECD (for Economic Co-operation and, OECD)). Further, AI fairness as equity (Holstein and Doroudi, 2019) is considered where the AI model assumes to be fair if it demonstrates an equal proportion of false positives and true positives for automatically predicted scores of different gender written responses. This section aims to dissect these aspects, providing a clearer understanding of the extent and implications of gender issues in AI systems.
Unraveling Myths: Misunderstanding Missing Information and flawed conclusions are almost always behind common myths about the gender bias of AI. One such myth is the belief that AI systems are naturally unbiased since they rely on algorithms and data. But as Bolukbasi et al. (2016) and Hardt et al. (2016) show, biases are inherent not only to the training data on which algorithmic design is based but even to its mechanisms of operation; they can result from seemingly gender-neutral words or expressions having a greater affinity with the Another myth is that the gender bias in AI lies wholly at a technical level, where it depends on solving these difficulties by gathering more data or writing better algorithms. These are essential factors, as recognized by Lu et al. (2020) and Manresa-Yee and Ramis (2021). However, overcoming gender bias in AI requires a more profound awareness of societal structures and cultural values.
Confronting Realities: Gender bias in AI is complex and multilayer, both quantitatively technical and qualitatively social. A type of bias is pseudo-AI Bias Zhai and Krajcik (2022) that reflects society’s stereotypes. Bias can appear in various forms, such as distorted dataset representations or biased interpretations of results from AI.
Impact on Education: In educational settings, this bias affects outcomes. It means unequal assessment, reinforced gender stereotypes, and barriers to genuine educational opportunity. Madaio et al. (2022) analyze structural injustices that may be propagated through AI applications in educational contexts. They call for more comprehensive approaches to understanding these biases.
Debunking Myths and Embracing Realities: The bottom line is that both in education and elsewhere, slaying myths about AI gender bias – along with being aware of it– can lead to developing fairer, more equitable systems for all learners. This requires an integrated, multi-faceted approach that incorporates the ideas of gender studies experts and people from practice in ethical AI design and educational theory to ensure our applications are not simply making existing dignities unequal. To understand and resolve AI gender discrimination in education, we must abandon myths and face up to the actual, much more complicated ways that various biases both exist within and arise from systems of artificial intelligence. Such an understanding is critical to creating AI applications that are fair and effective in different educational settings.
Study’s Approach to Bias, Disparities, and Fairness Analysis: The study employs a rigorous methodology to analyze gender issues in automatic scoring. We have fine-tuned standard BERT (similar to (Liu et al., 2023)) and GPT-3.5 (identical to (Latif and Zhai, 2023b)), using more than 1000 human expert’s rated student responses, proportionally distributed among male, female, and mixed datasets across six assessment items. The study measured bias by pair t-test of scoring accuracy between different gender-specifically trained models (Mowery, 2011); the gender disparities by mean score gaps by gender (Wang and Brown, 2007; Wilson et al., 2023); and fairness by EO (Hardt et al., 2016). These methods comprehensively show how gender issues manifest in AI scoring systems for mixed and separately trained models.
Gender-Neutral Nature of Mixed Trained Models: The study’s core hypothesis is that AI models trained using mixed-gender datasets don’t demonstrate gender prejudice in automatic scoring when tested on male and female responses. At its core is the notion that AI algorithms can only learn more gender-neutral patterns with a greater diversity of training data, thus eliminating biases resulting from unbalanced or skewed datasets.
Setting Ground to Avoid AI gender issues: A critical element of this study is its aim to demystify AI and gender problems. The study uses a data-driven, empirical approach to show that difficulties in AI surrounding gender are not immovable. Through careful and reasoned training methods, they can be solved. This goes against the prevailing misconception that all AI systems are biased or that arriving at bias-free AI is too complicated.
Implications for AI Development: These findings are significant for AI development in educational environments. They say that training data must be carefully thought about, and applying an appropriate analysis technique can result in fairer, more equitable AI systems. This is consistent with the broader debate on mainstreaming ethics in AI development, which stresses that all aspects of theory should be used to train AIs.
In short, this study adds to the ongoing efforts of many gender issues in AI through targeted training and rigorous analysis. It also provides empirical evidence and a conceptual framework, paving the way for adequate technical knowledge and full moral responsibility to develop responsible AI systems that do not fall prey to many wrong beliefs.
3. Methodology
This study aims to investigate gender issues in AI-based scoring systems based on state-of-the-art Language Models BERT and GPT-3.5 models. Large Language models tend to revolutionize education technology (Latif et al., 2023) and raise potential challenges, such as bias, disparity, and fairness, which must be addressed. A comprehensive methodology was adopted, comprising different training and testing phases for models and employing various statistical techniques to evaluate gender bias, disparities, and fairness. Figure 1 presents an overview of the research design.
3.1. Dataset
The study utilized a meticulously categorized dataset of student-written responses to six assessment items, each classified under the multi-class category. The datasets were categorized based on reported gender (i.e., male and female; we eliminated the data that did not report gender information), and a mixed dataset incorporated responses from both genders. Based on the availability of data, we used the maximum of available data while maintaining the ratio between training and testing data for each item by gender. Consequently, we randomly identified three training datasets for each item: mixed, male, and female training datasets and three testing datasets for each item. The detailed composition of each assessment item’s dataset is presented in Table 1.
This comprehensive dataset facilitated a nuanced analysis of gender bias in AI-based scoring systems, ensuring robust and broadly applicable study findings.
Table 1. Summary of Datasets and Their Samples
Datasets Category mixed Training samples mixed Testing samples Male Training Male Testing Female Training Female Testing
falling weights
Multi-class
1148
230
342
87
478
120
gelatin
Multi-class
918
230
344
86
477
120
bathtub
Multi-class
916
230
364
92
450
113
sandwater1
Multi-class
915
229
364
92
450
113
sandwater2
Multi-class
913
229
364
92
449
113
two boiling situations
Multi-class
909
228
363
91
448
113
3.2. Experimental Setup
The experimental setup involved rigorous training and testing using multiple training and testing datasets.
Training Phase: Each dataset undergoes a unique training cycle:
  • A mixed-trained model trained with the mixed training dataset including both genders.
  • A male-trained model trained exclusively with data from males.
  • A female-trained model trained solely with data from females.
Using this approach, we developed three distinct language models for each assessment item, fine-tuned using both BERT and GPT-3.5, respectively.
Testing Phase: Each trained model undergoes multiple testing phases to detect potential biases. Specifically,
  • Experiment 1. The model trained on the mixed dataset was tested against mixed testing data and separately against testing data from males and females. We used both BERT and GPT-3.5.
  • Experiment 2. Models trained on a single gender’s data were tested on their respective gender’s test data and against the test data from the opposite gender. Similarly, we used both BERT and GPT-3.5.
This experimental design aimed to rigorously assess any inherent biases in the models’ predictions across different datasets and gender divisions.
3.3. Accuracy Difference (Pair t-Test)
The accuracy difference between models was assessed using a paired t-test, a statistical method ideal for comparing the means of two related groups. Mowery (Mowery, 2011) describes the paired t-test as a powerful tool for detecting differences in paired observations. Our study involved comparing the accuracies of models trained on different datasets.
Mathematically, the paired t-test is expressed as:
(1) $t=\frac{\overline{D}}{s_{D}/\sqrt{n}}$
where $\overline{D}$ is the mean of the differences between paired observations, $s_{D}$ is the standard deviation of the differences, and $n$ is the number of pairs. A p-value greater than 0.05 indicated no significant difference in accuracies between the models, suggesting minimal gender bias.
3.4. Mean Score Gap
Mean Score Gap (MSG) was utilized to evaluate the disparity in scores between different genders. MSG can be generated either by human experts or machine algorithmic models. Given that human-assigned scores have undergone rigorous scrutiny, we used MSG generated by human-assigned scores as the standard to evaluate machine-generated MSG. It can be seen that if machine-generated MSG is larger than human-generated MSG for the same testing data, the machine exacerbates the gender difference and may potentially create inequity. Wang and Brown (2007) and Wilson et al. (2023)highlighted the importance of such comparisons in their study on automated essay scoring and automatic scoring of written scientific arguments. MSG is calculated as:
(2) $MSG=\frac{1}{n}\sum_{i=1}^{n}|score_{male,i}-score_{female,i}|$
where $score_{gender1,i}$ and $score_{gender2,i}$ are the scores assigned to the $i^{th}$ response by models trained on Male and Female datasets, respectively, and $n$ is the total number of responses. A threshold of MSG ¡ 0.2 was used to indicate acceptable score disparity.
3.5. Equalized Odds
Equalized Odds, as defined by Hardt et al. (Hardt et al., 2016), involves assessing the equality of true and false positive rates across groups. It is a crucial measure of fairness in predictive models. The Equalized Odds (EO) criterion is given by:
(3) $EO=max(|TPR_{male}-TPR_{female}|,|FPR_{male}-FPR_{female}|)$
where $TPR_{male}$ and $TPR_{female}$ are the true positive rates for Male and Female, respectively, and $FPR_{male}$ and $FPR_{female}$ are the false positive rates for the two genders. An EO value less than 0.01 indicates a fair model with minimal gender bias.
These three methods - Paired t-test, MSG, and Equalized Odds - provide a comprehensive approach to evaluating gender bias, disparities, and fairness in AI-based scoring systems, ensuring the ethical uses of AI models. The methodology encompassed a detailed experimental setup with rigorous training and testing phases and statistical analyses to assess gender issues in AI-based scoring systems comprehensively.
4. Results
4.1. Scoring Accuracy Difference Evaluation
A paired samples t-test was performed to evaluate whether there was a statistical difference between the scoring accuracy for male and female predicted scores before and after randomly mixing both genders’ responses.
4.1.1. BERT Model Results
[…]
4.1.2. GPT-3.5 Turbo Model Results
[…]
Interpretation: Both the BERT and GPT-3.5 models demonstrated consistent performance across mixed and gender-specific datasets, suggesting minimal gender biases.
Figure 2. Mean score gaps between male and female testing data: Comparing human-graded scores and fine-tuned LLMs’ scores: Mixed training Model (Left), Male trained Model (Center), and Female trained Model (Right) comparison line plots of BERT (Top) and GPT3.5 (Bottom) fine-tuned models’s scores.
4.2. Mean Score Gap
To verify our assumption, we designed two experiments to compare the difference of MSGs between human- and machine-graded scores. We tested the three models, Mixed-, male-, and female-trained models, and compared each of them with human-generated MSG.
4.2.1. Experiment 1: Mean Score Gap for the Mixed-Training Models
The first experiment analyzed the MSGs between human and mix-trained machine-graded scores for BERT and GPT-3.5, respectively. We tested the model performance with mixed, male, and female testing datasets, respectively.
For the BERT model, the MSG of the six tasks yielded from human-assigned scores on the mixed testing dataset was $M_{mixed}=0.032$, the male testing data was $M_{male}=0.056$, and the female testing data was $M_{female}=0.042$. Correspondingly, MSG generated by the mixed-trained BERT model on the three testing datasets were $M_{mixed}=0.022$, $M_{male}=0.048$, and $M_{female}=0.033$. The results indicate that the MSG between mixed test datasets against humans ($MD_{mixed}=0.011$) is comparatively lower than that of males ($MD_{male}=0.08$) and females ($MD_{female}=0.09$).
In contrast, for the GPT-3.5 model, the MSG of the six tasks yielded from human-assigned scores on the mixed testing dataset was $M_{mixed}=0.074$, the male testing data was $M_{male}=0.082$, and the female testing data was $M_{female}=0.112$. Correspondingly, MSG generated by the mixed-trained GPT-3.5 model on the three testing datasets were $M_{mixed}=0.072$, $M_{m}ale=0.072$, and $M_{female}=0.105$. The results indicate that the MSG between mixed test datasets against humans ($MD_{mixed}=0.002$) is comparatively lower than that of males ($MD_{male}=0.010$) and females ($MD_{female}=0.07$).
4.2.2. Experiment 2: Mean Score Gap for Gender Specifically Trained Models
To understand the MSGs generated by the gender-specifically-trained models, the second experiment compared the MSGs yielded by machine-graded scores and human-graded scores, respectively. For each assessment task, we calculated MSG on the male and female testing datasets, respectively. We first reported the BERT results, followed by the GPT-3.5 results.
For the BERT model, overall the two gender-specifically trained models demonstrated a larger average MSG for the six tasks, M = 0.107, compared to that generated by the human grader, M = 0.98. For example, ’sandwater2’, the MSGs were close for both testing datasets (Males: M=2.38, SD=0.67; Females: M=2.20, SD=0.73), akin to human MSGs (Males: M=2.33, SD=0.71; Females: M=2.20, SD=0.74), respectively. When a gender-specifically trained male model was tested on gender-counterpart testing data, we found that both Male and Female-trained Models showed increased MSGs on both testing datasets compared to human-generated MSGs. For example, the Male model in ’twoboilingsituation’ scored Female testing data higher (M=2.42, SD=0.59) than Female testing data, indicating an increased gender disparities. Conversely, when gender-specific models were tested on the same gender testing datasets, they exhibited decreased gender disparities (Detailed results can be seen in Fig. 2). An anomaly noted was the similarity in standard deviations between the Mixed Model and human scores, indicating consistent scoring variance.
Similarly, the GPT-3.5, overall, the fine-tuned male- and female-trained models exhibited a larger average MSG for six tasks, M = 0.115, compared to that graded by the human expert, M = 0.98. For instance, for the item ’twoboilingsituation’, the MSGs were close for both testing datasets (Males: M=2.40, SD=0.88; Females: M=2.15, SD=0.83), mirroring human scores (Males: M=2.37, SD=0.96; Females: M=2.19, SD=0.87), respectively. When a gender-specifically trained model was tested on gender-counter part testing data, we found that both Male and female-trained models showed increased MSGs on both testing datasets compared to human-generated MSGs. For instance, the male Model in ’sandwater1’ had a lower score for females (M=2.12, SD=0.76) than males. Across all items, the Mixed testing’s scores closely resembled human graders, suggesting reduced bias (Detailed results can be seen in Fig. 2).
Interpretation: These results indicate that training with a mixed dataset in both BERT and GPT-3.5 models has shown a minimal MSG compared to gender-specific training. This outcome validates our hypothesis that a balanced mixed training dataset have significant traces toward gender disparitiess, as reflected in the reduced MSGs in the mixed models. The findings underscore the importance of inclusive training datasets in developing AI models that yield more equitable scoring outcomes.
4.3. Equalized Odds Evaluation
The study evaluated the Equalized Odds (EO) (Hardt et al., 2016) for BERT and GPT-3.5 models to assess fairness in gender predictions.
4.3.1. BERT Model Equalized Odds
The BERT model demonstrated an EO of 0.042 on the mixed testing dataset for the mixed-trained model, indicating a lower disparity in prediction fairness. In contrast, the gender-specifically-trained models exhibited higher EO values on the mixed testing dataset, with the Male model showing an EO of 0.107 and the Female model having an EO of 0.074. These results suggest that the mixed-trained model is more equitable in its predictions across genders than gender-specific models.
4.3.2. GPT-3.5 Model Equalized Odds
For the GPT-3.5 model, the mixed-trained model’s EO was 0.061 on the mixed testing dataset, again reflecting a reduced gradient and, thus, greater fairness in predictions. The Male and Female trained models showed EOs on the same testing dataset of 0.076 and 0.074, respectively. While these values are closer to that of the mixed model than the BERT model, they indicate a higher disparity in for the gender-specific models.
Interpretation: The mixed-trained models for both BERT and GPT-3.5 have lower EO values, suggesting more equitable predictions and higher fairness compared to gender-specific models. This outcome supports the hypothesis that training with a balanced dataset can reduce the EO gradient, contributing to the development of fair AI models.
5. Discussion
The research shows that gender bias is absent from gender-unbalanced training datasets for AI-based scoring systems. The results provided evidence to refute the widely held belief that AI inevitably reinforces gender biases (Bolukbasi et al., 2016; Lu et al., 2020). Our findings, however, suggest that algorithms that are trained on data that is imbalanced in terms of gender could exacerbate existing gaps in gender. The promise of balanced training in addressing gender equity is demonstrated by the decreased EO gradients and reduced MSG in models trained on mixed datasets. This discovery is vital in light of Hardt et al. (2016), who highlight the significance of opportunity equality in supervised learning.
Our research indicates that balanced mixed training datasets have the potential to improve the equality of AI-based assessments in education, where the fairness of AI is crucial (Li et al., 2022). This strategy emphasizes the need for inclusive and equitable AI activities, which aligns with the ethical principles Slimi (2023) suggests for AI in higher education.
According to Zhai and Krajcik (2022) concept of pseudo AI bias, biases in AI might be deeply embedded and subtle rather than always obvious. This study’s demonstration of the balanced mixed training model’s gender-neutral behavior offers a path forward for resolving such unfounded allegations about AI biases. Gender stereotypes can be prevented from unintentionally being reinforced by AI models by guaranteeing varied representation in training data.
The literature supporting debiasing AI systems (Manresa-Yee and Ramis, 2021; Nadeem et al., 2022) aligns with our findings. This work shows that balanced training produces more equitable AI outcomes, which advances the larger attempts to develop impartial and fair AI models. The method offers a workable answer to this widespread problem by refuting that AI systems are bound to reproduce societal prejudices.
The study’s findings strongly support the gender issue of balanced mixed training in artificial intelligence for automated scoring. In addition to addressing the pressing issues of equality and fairness in AI-based scoring systems, our method offers proof against mechanical pseudo-bias in more general AI applications.
6. Conclusion
The analysis using BERT and GPT-3.5 on student written responses has significantly advanced our understanding of gender bias, disparities, and fairness in AI-based scoring systems. This study demonstrates that gender-unbalanced training has insignificant traces of gender bias. The findings challenge prevailing narratives about the inevitability of AI bias. We found that models trained on mixed-gender data, as opposed to gender-specific data, produced lower MSG when compared with human-generated results, implying that training data lacking gender balance might lead to algorithms exacerbating gender disparities. The EO analysis revealed that models trained on mixed-gender data achieved fairer outcomes than gender-specific ones. The results indicate that while gender-imbalanced data may not inherently lead to scoring bias, they can amplify gender disparities and diminish scoring fairness.
Acknowledgment
This study secondary analyzed data from a project supported by the Institute of Education Sciences (grant number R305A160219). The authors acknowledge the funding agencies and the project teams for making the data available for analysis. The project is partially supported by NSF (grant number 2101104). The findings, conclusions, or opinions herein represent the authors’ views and do not necessarily represent the views of personnel affiliated with the funding agencies.